OpenStudy (anonymous):

write the first four terms of the sequence for n=1,2,2 and 4 f(n)=6n^2

OpenStudy (anonymous):

you just put those numbers in. 6(1)^2, 6(2)^2, 6(3)^2, 6(4)^2, which would be 36, 144, 324, 576

OpenStudy (anonymous):

6(1)^2 = 6 6(2)^2= 24 6(3)^2= 54 6(4)^2= 96 elgen ur wrong

OpenStudy (anonymous):

6*1^2=6*1=6 6*2^2=6*4=24 6*3^2=6*6=36 6*4^2=6*6=48

OpenStudy (anonymous):

oops, that third one should have been 6*9=54

OpenStudy (anonymous):

its not being multiplied by two, its being squared

OpenStudy (anonymous):

ya so BOOM

OpenStudy (anonymous):

oh lol, for some reason i read it as if they were in a parenthesis. probably should pay more attention, sorry friends :P

OpenStudy (anonymous):

np, thanks for asking, now you know :D